ALGEBRAIC GEOMETRY SEMINAR IN JAN-APRIL 2004

The seminar is on each MONDAY at 13.45 hours.

ROOM : 217, NEW BUILDING, IMSC.

(Seminar schedule in Sept-Dec 2003 is on p.2).

—————————————–SCHEDULE—————————————–

DATE:—-:SPEAKER :————————–TITLE—————- 05.01.04: C.S.Dalawat: The Chow group of a rational surface over a local field. (Time: 11.30 hours 19.01.04: Arijit Dey : Kodaira-Spencer’s theorem on deformations. (Time: 13.45 a.m) 19.01.04 : S.Ramanan : Moduli spaces as Fano varieties. (Time: 15.30) 20.01.04 : Ulf Persson : Geometry of threefolds (Time: 14.00)

DATE :— :SPEAKER : ———————— TITLE ————————————

15.09.03: Jaya Iyer : Gromov Witten invariants 1 (Introduction). 22.09.03: Jaya Iyer : Gromov Witten invariants 2 (Axiomatic approach).

29.09.03: Kapil Paranjape: Enumerative Geometry. 06.10.03 : Kapil Paranjape: Continued. 13.10.03 : Kapil Paranjape: About M̄(X, β).

20.10.03 : D.S.Nagaraj: The construction of M̄_{0,n}(P^{r}, d). 27.10.03 : D.S.Nagaraj: M̄_{o,n}(X, β) and its boundary. 03.11.03 : D.S.Nagaraj: Continued.

10.11.03 : P.Sankaran: The small quantum cohomology of the Grassmanian 1. 17.11.03 : P.Sankaran: The small quantum cohomology of the Grassmanian 2.

24.11.03 : S.Ramanan : On Mori’s Theorem. 01.12.03 : S.Ramanan : Continued.

01.12.03 : D.N.Verma : Study of higher syzygies of some Grassmanians. (Time: 11.30). 23.12.03 : Vijaylaxmi Trivedi : Semi-stability and Hilbert-Kunz multiplicities. 23.12.03 : V.Srinivas : A problem in Linear Algebra. (Time: 15.30)